# Not All Rollouts are Useful: Down-Sampling Rollouts in LLM Reinforcement Learning

### v3

@@ 4.2 Effect of Rollout and Update Sizes $(n,m)$
A key practical question for PODS adoption is how to choose the rollout size $(n)$ and training batch size $(m)$. While a larger $n$ provides more diverse rollouts for selection, it also increases inference costs. Meanwhile, a smaller $m$ reduces update costs but may provide insufficient training signal. As shown in Fig.˜4, we systematically study these trade-offs to provide deployment guidance.
@@
These results suggest down-sampling ratio of $2$ to $4$ (i.e., $m=16$ and $n=32$ or $64$ in Fig.˜4(a)) provides an effective balance of performance and efficiency. For resource-constrained settings, aggressive down-sampling ratios up to $16$ remain viable, while memory-rich environments can benefit from larger rollout pools up to hardware limits.
@@ 5 Conclusion and Discussion
We introduced PODS—a lightweight, algorithm-agnostic framework that addresses a fundamental bottleneck in modern RLVR training: the asymmetry between embarrassingly parallel rollout generation and memory-intensive policy updates. PODS generates large batches of rollouts in parallel and updates the policy on only an informative subset chosen by the max-variance rule. Our analysis shows that the optimal subset can be found in $O(n\log n)$ time. This simple yet principled approach consistently outperforms standard GRPO under equal wall-clock budgets, delivers at least a $1.7\times$ speedup and reaching higher final accuracy across diverse model architectures, scales, and deployment scenarios. Our ablation study shows that the performance of PODS is robust over a wide range of down-sampling ratios provided $m$ is not too small, empirically confirming our method’s efficacy.
@@
Our evaluation focuses on mathematical-reasoning tasks with verifiable rewards. Other domains such as open-ended dialogue or code generation may exhibit distinct dynamics of the algorithms. Moreover, in workloads that demand greater prompt diversity, similar gains might be obtained by processing more prompts per iteration with fewer rollouts per prompt and accumulating gradients across prompts—an alternative path to address the inference-update asymmetry. Finally, because PODS alters the training rollout distribution through selective down-sampling, it behaves off-policy and may be unsuitable when strict on-policy guarantees are required.
Future work.
The algorithm-agnostic nature of PODS enables integration with value-based methods like PPO and emerging RL approaches. Exploring whether PODS can enhance state-of-the-art model performance represents a pzromising research problem. Additionally, investigating adaptive down-sampling strategies that evolve throughout training could further optimize the learning dynamics. Exploring theoretically principled approaches to balance the trade-off between prompt diversity and rollout quantity per prompt also warrants investigation.
@@ A.1 Reward Functions
The specific reward functions we use in our experiments are listed below.
Accuracy ($\mathbf{1.0}$ for correct, $\mathbf{0.0}$ for incorrect): Mathematical correctness using parsing and symbolic verification. The reward function extracts mathematical expressions from both the model’s response and ground truth solution, then uses symbolic verification to determine equivalence.
Format ($\mathbf{1.0}$ for compliant, $\mathbf{0.0}$ for non-compliant): Structured response formatting that requires reasoning to be enclosed within <think> tags and the final answer within <answer> tags, following the exact pattern $\n...\n</think>\n$$\n...\n</answer>$.
@@ A.2 Hyperparameters
Setting (a) (b) (c) (d) (e)
Optimizer AdamW AdamW AdamW AdamW AdamW
Max Sequence Length $1024$ $1024$ $1024$ $2048$ $2048$
Lora Rank $64$ $64$ $64$ N/A N/A
Lora Alpha $64$ $64$ $64$ N/A N/A
KL Coefficient $0.00$ $0.00$ $0.04$ $0.00$ $0.00$
Learning Rate $5\times 10^{-6}$ $5\times 10^{-6}$ $2\times 10^{-6}$ $2\times 10^{-5}$ $1.5\times 10^{-5}$
Weight Decay $0.1$ $0.1$ $0.1$ $0.1$ $0.1$
Grad Clipping $1.0$ $1.0$ $1.0$ $1.0$ $1.0$
GA Steps (GRPO-PODS) 11 11 11 44 44
Rollout Batch Size (GRPO-PODS) $64$ $32$ $64$ $128$ $128$
Update Batch Size (GRPO-PODS) $16$ 88 $16$ $32$ $32$
Effective $n$ (GRPO-PODS) $64$ $32$ $64$ $512$ $512$
Effective $m$ (GRPO-PODS) $16$ 88 $16$ $128$ $128$
Down-Sampling Ratio 44 44 44 44 44
GA Steps (GRPO) 11 11 11 N/A N/A
Rollout Batch Size (GRPO) $16$ 88 $16$ N/A N/A
Update Batch Size (GRPO) $16$ 88 $16$ N/A N/A
Effective $n$ (GRPO) $16$ 88 $16$ N/A N/A
Effective $m$ (GRPO) $16$ 88 $16$ N/A N/A
GA Steps (GRPO-GA) N/A N/A N/A $16$ $16$
Rollout Batch Size (GRPO-GA) N/A N/A N/A $32$ $32$
Update Batch Size (GRPO-GA) N/A N/A N/A $32$ $32$
Effective $n$ (GRPO-GA) N/A N/A N/A $512$ $512$
Effective $m$ (GRPO-GA) N/A N/A N/A $512$ $512$
@@
For experiment settings (d) and (e), we ensured a fair comparison between GRPO-PODS and GRPO-GA by matching the total number of rollouts (effective $n$) generated per prompt. This was done by equating the product of rollout batch size and GA steps across both methods. In the open-r1 implementation, GA steps determine both rollout generation and training updates. For example, with rollout batch size $128$ and GA steps $4$, the effective $n$ is $128\times 4=512$. We fixed this number at $512$ for both GRPO-PODS and GRPO-GA.
For GRPO-PODS, each rollout batch was down-sampled by a factor of $4$, resulting in an update batch size of $32$ and an effective $m$ of $32\times 4=128$. Because down-sampling is applied directly after generating each batch rather than after aggregation, GRPO-GA must increase GA steps by $4\times$ (from $4$ to $16$) to maintain the same effective $n$. This adjustment ensures that both variants process an equal number of rollouts while respecting their structural differences.
A.3 Comparing Different Down-Sampling Rules
We study the effect of different down-sampling rules on the performance of GRPO-PODS in this section. We conduct experiments on the GSM8K dataset with one L40S GPU. We set the rollout size $n=64$ and the update size $m=16$, and we compare three different down-sampling rules: (1) max-variance down-sampling, (2) max-reward down-sampling, and (3) random down-sampling. The results are shown in Fig.˜5. We observe that the max-variance down-sampling rule consistently outperforms both the max-reward and random down-sampling rules across all settings. This indicates that the max-variance down-sampling rule is effective in selecting informative rollouts for training.
Figure 5: Performance and per-step run time comparison of GRPO-PODS with the max-variance, max-reward and random down-sampling rules. The training is conducted on the GSM8K dataset with one L40S. For the performance comparison, the x-axis shows the training time, and the y-axis shows the accuracy on the test set. The shaded area represents 1.96 times the standard error of the mean.
@@ A.4 PODS’ Speed Up Ratio Over GRPO
In Fig.˜3, we observe that GRPO-PODS consistently outperforms GRPO in terms of performance as the training proceeds. For each of the five plots in Fig.˜3, we compute the speed up ratio of GRPO-PODS over GRPO, i.e., the ratio between the time taken by GRPO and that taken by GRPO-PODS to reach $0.99\times$ the peak performance of GRPO. The results are shown in Table˜3. We observe that our method achieves a speed up ratio between $1.7\times$ and $3.0\times$ over GRPO across the settings.
@@
Setting (a) (b) (c) (d) (e)
Speed Up Ratio $\times 2.0×2.0\times$ $3.0\times$ $\times 1.7×1.7\times$
A.5 Average Completion Length Over Time
We include additional evaluation of the average completion length over the training time for each of the experiments we conduct in Section˜4. We present the average completion length results in Figs.˜6, 7 and 8, in correspondence to Figs.˜3, 4 and 5 respectively. In most of the cases, we observe that the average completion length stays relatively stable over the training time.
(a) Training Qwen2.5 (3B) on GSM8K with one L40S GPU
(b) Training Qwen2.5 (3B) on MATH with one L40S GPU
(c) Training Llama3.2 (3B) on GSM8K with one L40S GPU
(d) Training Qwen2.5 (3B) on GSM8K with 8 H100 GPUs
(e) Training Qwen2.5 (7B) on GSM8K with 8 A100 GPUs
@@ A.5 Average Completion Length Over Time
Figure 8: Average completion length over time of the trained models in Section˜A.3’s experiments. The x-axis shows the training time, and the y-axis shows the average completion length in tokens. The shaded area represents 1.96 times the standard error of the mean.

### v4

@@ 4.2 Effect of Rollout and Update Sizes $(n,m)$
A key practical question for PODS adoption is how to choose the rollout size $(n)$ and training batch size $(m)$. While a larger $n$ provides more diverse rollouts for selection, it also increases inference costs. Meanwhile, a smaller $m$ reduces update costs but may provide insufficient training signal. As shown in Fig.˜4, we study these trade-offs on experimental setting (a) to provide deployment guidance.
@@
These results suggest down-sampling ratio of $2$ to $4$ provides an effective balance of performance and efficiency. For resource-constrained settings, aggressive down-sampling ratios up to $16$ remain viable, while memory-rich environments can benefit from larger rollout pools up to hardware limits.
4.3 Comparing Different Down-Sampling Rules
We study the effect of different down-sampling rules on the performance of GRPO-PODS in this section. We conduct experiments on experimental setting (a). We compare four different down-sampling rules: (1) max-variance down-sampling, (2) max-reward down-sampling, (3) random down-sampling, and (4) percentile down-sampling. The results are shown in Fig.˜5. We observe that the max-variance down-sampling rule consistently outperforms the other down-sampling rules throughout the training. This indicates that the max-variance down-sampling rule is effective in selecting informative rollouts for training.
Figure 5: Performance of GRPO-PODS with the max-variance, max-reward, random and percentile down-sampling rules. The training is conducted on the GSM8K dataset with one L40S. The x-axis shows the training time, and the y-axis shows the accuracy on the test set. The shaded area represents 1.96 times the standard error of the mean. The right panel shows a Zoomed-in view of the left one.
@@ 5 Conclusion and Discussion
We introduced PODS, a lightweight framework that addresses a fundamental bottleneck in modern RLVR training: the asymmetry between embarrassingly parallel rollout generation and memory-intensive policy updates. PODS generates large batches of rollouts in parallel and updates the policy on only an informative subset chosen by the max-variance rule. Our analysis shows that the optimal subset can be found in $O(n\log n)$ time. This simple yet principled approach consistently outperforms standard GRPO under equal wall-clock budgets, delivers at least a $1.7\times$ speedup and reaching higher final accuracy across diverse model architectures, scales, and deployment scenarios. Our ablation study shows that the performance of PODS is robust over a wide range of down-sampling ratios provided $m$ is not too small, empirically confirming our method’s efficacy.
@@
As we emphasize in the paper, our evaluation focuses on RLVR tasks where response correctness is verifiable. Other domains such as open-ended dialogue may exhibit distinct dynamics of the algorithms. Moreover, in workloads that demand greater prompt diversity, similar gains might be obtained by processing more prompts per iteration with fewer rollouts per prompt and accumulating gradients across prompts, which is an alternative path to address the inference-update asymmetry. Finally, because PODS alters the training rollout distribution through selective down-sampling, it behaves off-policy and may be unsuitable when strict on-policy guarantees are required despite its improved empirical performance.
Scope of claims. Our empirical and theoretical claims in this paper are specific to PODS instantiated within GRPO. While we expect the inference-update asymmetry to be present in other RLVR pipelines, demonstrating that down-sampling yields comparable benefits for other RL methods (e.g., PPO-style objectives or value-based variants) requires separate analysis and experimentation.
Discussion and future work.
PODS is a general framework that admits different down-sampling rules. In this paper, we focus on the max-variance rule due to its superior empirical performance in common RLVR settings. In some scenarios, other rules might be more effective. For example, when different prompts cause the model to have highly varying reward distributions, applying the max-variance rule across all rollouts may lead to over-sampling from a small subset of prompts with extreme difficulty levels. In such cases, applying the max-variance rule within each prompt and then selecting a balanced subset across prompts may be more effective. More broadly, the down-sampling rule can be designed to take into account more information beyond the reward values, such as the rollouts’ entropy, similarity measures between rollouts, or even a target reward distribution that we wish to down-sample towards. Exploring these alternative down-sampling rules and their empirical performance across different settings is an interesting direction for future work.
Reproducibility statement.
We describe the rewards used in Section˜A.1, list the key hyperparameters in Section˜A.2, and an anonymized version of the code used to run the experiments in this paper is attached to our submission as supplementary material on OpenReview. We will publicly release the code on GitHub, and include its link in the next version of our paper. We note that all of the datasets used in this paper are open-source, and the models we use are all open-weight and available publicly on HuggingFace.
@@ A.1 Reward Functions
We list the reward functions we use in our experiments below.
Accuracy ($\mathbf{1}$ for correct, $\mathbf{0}$ for incorrect): For mathematical reasoning tasks, we use a binary reward indicating whether the final answer is correct. This is computed by comparing the model’s final answer to the ground truth solution using parsing and symbolic verification, which allows for robust equivalence checking even when the model’s answer is not in the exact same format as the ground truth. For Chemistry problems, the answer is always a letter in {A, B, C, D}, so we can directly compare with the correct answer.
Format ($\mathbf{1}$ for compliant, $\mathbf{0}$ for non-compliant): For a response to be considered format-compliant, it must follow the exact XML structure we specify for reasoning and answer. This requires reasoning to be enclosed within <think> tags and the final answer within <answer> tags, following the exact pattern $\n...\n</think>\n$$\n...\n</answer>$.
@@ A.2 Hyperparameters
Setting (a) (b) (c) (d) (e) (f)
Optimizer AdamW AdamW AdamW AdamW AdamW AdamW
Max Sequence Length $1024$ $1024$ $1024$ $1024$ $2048$ $2048$
Lora Rank $64$ $64$ $64$ $64$ N/A N/A
Lora Alpha $64$ $64$ $64$ $64$ N/A N/A
KL Coefficient $0.00$ $0.04$ $0.00$ $0.00$ $0.00$ $0.00$
Learning Rate $5\cdot 10^{-6}$ $2\cdot 10^{-6}$ $5\cdot 10^{-6}$ $5\cdot 10^{-6}$ $2\cdot 10^{-5}$ $1.5\cdot 10^{-5}$
Weight Decay $0.1$ $0.1$ $0.1$ $0.1$ $0.1$ $0.1$
Grad Clipping $1.0$ $1.0$ $1.0$ $1.0$ $1.0$ $1.0$
GA Steps (GRPO-PODS) 11 11 11 11 44 44
Rollout Batch Size (GRPO-PODS) $64$ $64$ $32$ $64$ $128$ $128$
Update Batch Size (GRPO-PODS) $16$ $16$ 88 $16$ $32$ $32$
Effective $n$ (GRPO-PODS) $64$ $64$ $32$ $64$ $512$ $512$
Effective $m$ (GRPO-PODS) $16$ $16$ 88 $16$ $128$ $128$
Down-Sampling Ratio 44 44 44 44 44 44
GA Steps (GRPO) 11 11 11 11 N/A N/A
Rollout Batch Size (GRPO) $16$ 88 $16$ $16$ N/A N/A
Update Batch Size (GRPO) $16$ 88 $16$ $16$ N/A N/A
Effective $n$ (GRPO) $16$ 88 $16$ $16$ N/A N/A
Effective $m$ (GRPO) $16$ 88 $16$ $16$ N/A N/A
GA Steps (GRPO-GA) N/A N/A N/A N/A $16$ $16$
Rollout Batch Size (GRPO-GA) N/A N/A N/A N/A $32$ $32$
Update Batch Size (GRPO-GA) N/A N/A N/A N/A $32$ $32$
Effective $n$ (GRPO-GA) N/A N/A N/A N/A $512$ $512$
Effective $m$ (GRPO-GA) N/A N/A N/A N/A $512$ $512$
@@
For experiment settings (e) and (f), we ensure a fair comparison between GRPO-PODS and GRPO-GA by matching the total number of rollouts (effective $n$) generated per prompt. This is done by equating the product of rollout batch size and GA steps across both methods. In the open-r1 implementation, GA steps determine both rollout generation and training updates. For example, with rollout batch size $128$ and GA steps $4$, the effective $n$ is $128\times 4=512$. We fixed this number at $512$ for both GRPO-PODS and GRPO-GA. For GRPO-PODS, each rollout batch is down-sampled by a factor of $4$, resulting in an update batch size of $32$ and an effective $m$ of $32\times 4=128$. Because down-sampling is applied directly after generating each batch rather than after aggregation, GRPO-GA must increase GA steps by $4\times$ (from $4$ to $16$) to maintain the same effective $n$. This adjustment ensures that both variants process an equal number of rollouts while respecting their structural differences.
A.3 On Advantage Normalization
In Section˜3.2, we mention that in GRPO-PODS, we first apply the down-sampling to the rollout batch, and then compute the advantage normalization statistics (mean and standard deviation) using the down-sampled batch. This design ensures the advantage normalization is consistent with the actual data used for policy updates, so that each update batch has a total advantage of 0. In this section, we study the effect of this design choice by comparing it with an alternative design where the advantage normalization statistics are computed using the full rollout batch before down-sampling. We conduct this comparison in the experimental setting (a) where we train Qwen2.5 (3B) on GSM8K with one L40S GPU. The results are shown in Fig.˜6. We observe that computing the advantage normalization statistics using the down-sampled batch leads to better performance than using the full rollout batch. We hypothesize that the fact that the advantage normalization is consistent with the actual data used for policy updates is essential to the improved performance.
Figure 6: Performance and per-step run time comparison of the two design choices of advantage normalization in GRPO-PODS with max-variance down-sampling in experimental setting (a). “After” refers to computing the advantage normalization statistics using the down-sampled batch, while “Before” refers to doing so using the full rollout batch before down-sampling. The x-axis shows the training time, and the y-axis shows the accuracy on the test set. The shaded area represents 1.96 times the standard error of the mean.
@@ A.4 PODS’ Speed Up Ratio Over GRPO
In Fig.˜3, we observe that GRPO-PODS consistently outperforms GRPO in terms of performance as the training proceeds. For each of the six plots in Fig.˜3, we compute the speed up ratio of GRPO-PODS over GRPO, i.e., the ratio between the time taken by GRPO and that taken by GRPO-PODS to reach $0.99\times$ the peak performance of GRPO. The results are shown in Table˜3. We observe that our method achieves a speed up ratio between $1.7\times$ and $3.0\times$ over GRPO across the settings.
@@
Setting (a) (b) (c) (d) (e) (f)
Speed Up Ratio $2.0\times$ $3.0\times$ $2.0\times$ $1.8\times$ $\times 1.7×1.7\times$
A.5 Additional Evaluation for Experimental Settings (a) and (b)
To demonstrate that the performance improvement of GRPO-PODS over GRPO is not limited to the specific test set used in Fig.˜3, we evaluate the trained models on additional test sets for experimental settings (a) and (b). For both settings, we evaluate the trained models on the GSM8K Platinum test set, which is a contamination-resistant subset of GSM8K, and MATH. The results are shown in Fig.˜7. We observe that GRPO-PODS consistently outperforms GRPO across all the test sets, demonstrating the robustness of the performance improvement, and indicating the improvement is not due to overfitting to a specific test set.
(a) Training Qwen2.5 (3B) on GSM8K with one L40S GPU and Evaluating on GSM8K Platinum
(b) Training Qwen2.5 (3B) GSM8K with one L40S GPU and Evaluating on MATH
(c) Training Llama3.2 (3B) on GSM8K with one L40S GPU and Evaluating on GSM8K Platinum
(d) Training Llama3.2 (3B) on GSM8K with one L40S GPU and Evaluating on MATH
Figure 7: Additional performance comparison of standard GRPO and GRPO-PODS with max-variance down-sampling in experimental settings (a-b). The x-axis shows the training time, and the y-axis shows the accuracy on the test set. The shaded area represents 1.96 times the standard error of the mean.
A.6 Average Completion Length Over Time
We include additional evaluation of the average completion length over the training time for each of the experiments we conduct in Section˜4. We present the average completion length results in Figs.˜10, 8 and 9, in correspondence to Figs.˜3, 4 and 5 respectively. In most of the experimental settings, we observe that the average completion length stays relatively stable over the training time.
@@ A.6 Average Completion Length Over Time
Figure 9: Average completion length over time of the trained models in Section˜4.3’s experiments. The x-axis shows the training time, and the y-axis shows the average completion length in tokens. The shaded area represents 1.96 times the standard error of the mean.
(a) Training Qwen2.5 (3B) on GSM8K with one L40S GPU
(b) Training Llama3.2 (3B) on GSM8K with one L40S GPU
(c) Training Qwen2.5 (3B) on MATH with one L40S GPU
(d) Training Qwen2.5 (3B) on Chemistry with one L40S GPU
(e) Training Qwen2.5 (3B) on GSM8K with 8 H100 GPUs
(f) Training Qwen2.5 (7B) on GSM8K with 8 A100 GPUs